OpenStudy (anonymous):

solve: 5/9x-17=25-2/9x I got: 24

OpenStudy (anonymous):

(5/9)x-17=25-(2/9)x (7/9)x=42 x=54

OpenStudy (anonymous):

Unless the x's were actually part of the denominator

OpenStudy (luigi0210):

Yup she's right.. again :P She beat me to the answer

OpenStudy (anonymous):

It's only algebra, I'm sure you know everything too, and just happened to read the problem after me =P

OpenStudy (anonymous):

I see what I did wrong. just a careless mistake. thanks for the verification.:P

OpenStudy (anonymous):

Yup no problem! Happens to the best of us.